# JointFM-0.1: A Foundation Model for Multi-Target Joint Distributional Prediction

### 14 Mar 2026 v1

@@ Abstract
Despite the rapid advancements in Artificial Intelligence (AI), Stochastic Differential Equations (SDEs) remain the gold-standard formalism for modeling systems under uncertainty. However, applying SDEs in practice is fraught with challenges: modeling risk is high, calibration is often brittle, and high-fidelity simulations are computationally expensive. This technical report introduces JointFM, a foundation model that inverts this paradigm. Instead of fitting SDEs to data, we sample an infinite stream of synthetic SDEs to train a generic model to predict future joint probability distributions directly. This approach establishes JointFM as the first foundation model for distributional predictions of coupled time series — requiring no task-specific calibration or finetuning. Despite operating in a purely zero-shot setting, JointFM reduces the energy loss by 14.2% relative to the strongest baseline when recovering oracle joint distributions generated by unseen synthetic SDEs.
@@ 1 Introduction
Figure 1: JointFM, the digital quant. Traditional quantitative modeling follows a three-stage select-calibrate-simulate pipeline (top): choose a system of stochastic processes, fit parameters to historical data, and then simulate future paths [kloeden1992numerical, glasserman2003monte]. JointFM replaces this workflow by pretraining on a universe of synthetic SDE dynamics and directly predicting future joint probability distributions from context in a single forward pass (bottom).
@@
The Solution: A “Distributional Foundation Model” that predicts the full joint distribution of future outcomes directly from context (zero-shot) whenever a risk decision needs to be made [^1] We use the term “foundation model” not in the classical sense of a model that serves as a foundation but must be adapted to a specific task [bommasani2022opportunitiesrisksfoundationmodels] but for a model that serves as a generic, pretrained estimator for a vast collection of tasks. However, having a foundation model does not preclude further task-specific finetuning if desired., for instance, it becomes feasible to access the risk of a fund, like an ETF, based on its constituents on-the-fly, without needing to fit a complex multivariate SDE.
@@
  • Synthetic Physics Pretraining: Prior work has explored synthetic pretraining for zero-shot forecasting—TempoPFN [moroshan2026tempopfn] trains on a single univariate regime-switching Ornstein–Uhlenbeck SDE [uhlenbeck1930theory]. We scale this principle to a curriculum-controlled, infinite stream of diverse multivariate SDE systems—spanning correlated diffusions, jump processes, regime switching, and non-Markovian memory(though memory dynamics are not yet included in our training experiments).
@@ 2.1 Time-Series Foundation Models (TSFMs)
Recent advances in Time-Series Foundation Models (e.g., Chronos-2 [ansari2025chronos2], Moirai-2.0 [liu2026moirai20], TimesFM-2.5 [das2024decoderonly]) have demonstrated impressive zero-shot forecasting capabilities by leveraging large-scale pretraining on diverse datasets. However, a critical limitation of these models is their focus on univariate forecasting. They typically process each time series independently and do not predict the full joint probability distribution across targets. This is a crucial distinction: risk-aware decision-making under uncertainty—e.g., portfolio optimization, supply-chain hedging, or safety-critical control [mesbah2016stochastic]—requires coherent probabilistic scenarios that capture the cross-variable dependency structure, including tail dependencies and nonlinear co-movements. Predicting each target’s distribution in isolation, even accurately, is insufficient when the correlation structure is as important as the marginal distributions [embrechts2002correlation].
@@ 4 Full global correlation $\rho_{\text{global}}\in\mathbb{R}^{(M+N)\times(M+N)}$
5 Compound Poisson jumps [merton1976optionpricing_sciencedirect] $J(t,X_{t})\,dN_{t}$
@@ 3.3 Architecture Fundamentals & Distributional Heads
  • Skewed Student-t Mixture: Predicts weights, degrees of freedom ($\nu$), location ($\mu$), scale matrix ($\Sigma$), and skewness ($\alpha$) for a mixture of multivariate skewed Student-t components. This parametric form explicitly captures heavy tails and asymmetry, essential for financial risk modeling [azzalini2003distributions, mcculloch1989multivariateskewt].
@@ 4 Experiment: Zero-Shot Synthetic Distribution Recovery
Setup: We draw 200 unseen SDE systems at curriculum level 7 (the highest Markovian level, which includes all dynamics up to logistic regime switching, including jumps). Each system comprises $N=10$ coupled target series with no exogenous features ($M=0$). From each sampled SDE we simulate a single history realization of $T=504$ input steps (spanning a simulated time window of $t_{\text{in}}=2.0$), which serves as the model’s context. For the ground truth, we branch 1,000 independent future paths of $H=63$ steps ($t_{\text{out}}=0.25$) from the terminal state of the history, yielding a Monte Carlo approximation to the true conditional joint distribution $P(X_{T+1:T+H}\mid X_{1:T})$. All models produce 1,000 forecast samples, and metrics are computed independently at each of the 63 forecast horizons. JointFM is evaluated in pure zero-shot mode (no finetuning on the test instances); baselines are fitted directly to the observed history window.
@@
    DCC-GARCH: A classic multivariate GARCH model with dynamic conditional correlations, fitted to the history [engle2002dynamic].
@@
Results (Figure 2): We observe that JointFM consistently outperforms the baselines across all three metrics. When predicting 63 steps into the future, JointFM - Student-t achieves 10.5% lower energy loss, 12.7% lower marginal energy loss, and 10.1% lower CRPS-sum relative to the best-performing baseline. Averaged across all forecast horizons, JointFM - Student-t achieves 14.2% lower energy loss, 12.4% lower marginal energy loss, and 7.2% lower CRPS-sum relative to the best-performing baseline.

### 03 Apr 2026 v2

@@ Abstract
Despite the rapid advancements in Artificial Intelligence (AI), Stochastic Differential Equations (SDEs) remain the gold-standard formalism for modeling systems under uncertainty. However, applying SDEs in practice is fraught with challenges: modeling risk is high, calibration is often brittle, and high-fidelity simulations are computationally expensive. This technical report introduces JointFM, a foundation model that inverts this paradigm. Instead of fitting SDEs to data, we sample an infinite stream of synthetic SDEs to train a generic model to predict future joint probability distributions directly. This approach establishes JointFM as the first foundation model for distributional predictions of coupled time series — requiring no task-specific calibration or finetuning. Despite operating in a purely zero-shot setting, JointFM reduces the energy loss by 21.1% relative to the strongest baseline when recovering oracle joint distributions generated by unseen synthetic SDEs.
@@ 1 Introduction
Figure 1: JointFM, the digital quant. Traditional quantitative modeling follows a three-stage select-calibrate-simulate pipeline (top): choose a system of stochastic processes, fit parameters to historical data, and then simulate future paths [kloeden1992numerical, glasserman2003monte]. JointFM replaces this workflow by pretraining on a universe of synthetic SDE dynamics and directly predicting future joint probability distributions from context in a single forward pass (bottom). Image generated with Gemini.
@@
The Solution: A “Distributional Foundation Model” that predicts the full joint distribution of future outcomes directly from context (zero-shot) whenever a risk decision needs to be made [^1] We use the term “foundation model” not in the classical sense of a model that serves as a foundation but must be adapted to a specific task [bommasani2021opportunitiesrisksfoundationmodels] but for a model that serves as a generic, pretrained estimator for a vast collection of tasks. However, having a foundation model does not preclude further task-specific finetuning if desired., for instance, it becomes feasible to access the risk of a fund, like an ETF, based on its constituents on-the-fly, without needing to fit a complex multivariate SDE.
@@
  • Synthetic Physics Pretraining: Prior work has explored synthetic pretraining for zero-shot forecasting—TempoPFN [moroshan2025tempopfn] trains on a single univariate regime-switching Ornstein–Uhlenbeck SDE [uhlenbeck1930theory]. We scale this principle to a curriculum-controlled, infinite stream of diverse multivariate SDE systems—spanning correlated diffusions, jump processes, regime switching, and non-Markovian memory(though memory dynamics are not yet included in our training experiments).
@@ 2.1 Time-Series Foundation Models (TSFMs)
Recent advances in Time-Series Foundation Models (e.g., Chronos-2 [ansari2025chronos2], Moirai-2.0 [liu2025moirai20], TimesFM-2.5 [das2023decoderonly]) have demonstrated impressive zero-shot forecasting capabilities by leveraging large-scale pretraining on diverse datasets. However, a critical limitation of these models is their focus on univariate forecasting. They typically process each time series independently and do not predict the full joint probability distribution across targets. This is a crucial distinction: risk-aware decision-making under uncertainty—e.g., portfolio optimization, supply-chain hedging, or safety-critical control [mesbah2016stochastic]—requires coherent probabilistic scenarios that capture the cross-variable dependency structure, including tail dependencies and nonlinear co-movements. Predicting each target’s distribution in isolation, even accurately, is insufficient when the correlation structure is as important as the marginal distributions [embrechts2002correlation].
@@ 4 Full global correlation $\rho_{\text{global}}\in\mathbb{R}^{(M+N)\times(M+N)}$
5 Compound Poisson jumps [merton1976optionpricing] $J(t,X_{t})\,dN_{t}$
@@ 3.3 Architecture Fundamentals & Distributional Heads
  • Skewed Student-t Mixture: Predicts weights, degrees of freedom ($\nu$), location ($\mu$), scale matrix ($\Sigma$), and skewness ($\alpha$) for a mixture of multivariate skewed Student-t components. This parametric form explicitly captures heavy tails and asymmetry, essential for financial risk modeling [azzalini2003distributions].
@@ 4 Experiment: Zero-Shot Synthetic Distribution Recovery
Setup: We draw 1,000 unseen SDE systems at curriculum level 7 (the highest Markovian level, which includes all dynamics up to logistic regime switching, including jumps). Each system comprises $N=10$ coupled target series with no exogenous features ($M=0$). From each sampled SDE we simulate a single history realization of $T=504$ input steps (spanning a simulated time window of $t_{\text{in}}=2.0$), which serves as the model’s context. For the ground truth, we branch 1,000 independent future paths of $H=63$ steps ($t_{\text{out}}=0.25$) from the terminal state of the history, yielding a Monte Carlo approximation to the true conditional joint distribution $P(X_{T+1:T+H}\mid X_{1:T})$. All models produce 1,000 forecast samples, and metrics are computed independently at each of the 63 forecast horizons. JointFM is evaluated in pure zero-shot mode (no finetuning on the test instances); baselines are fitted directly to the observed history window.
@@
    DCC-GARCH: A classic multivariate GARCH model with dynamic conditional correlations, fitted to the history [engle2002dynamic]. We use the open-source implementation from https://github.com/srivastavaprashant/mgarch with a default multivariate normal distribution and patch scipy.optimize.minimize to enforcemaxiter=500.
@@
Results (Figure 2): We observe that JointFM consistently outperforms the baselines across all three metrics. When predicting 63 steps into the future, JointFM - Student-t achieves 19.6% lower energy loss, 15.9% lower marginal energy loss, and 9.1% lower CRPS-sum relative to the best-performing baseline. Averaged across all forecast horizons, JointFM - Student-t achieves 21.1% lower energy loss, 15.8% lower marginal energy loss, and 7.3% lower CRPS-sum relative to the best-performing baseline.